One number is 38 less than another. Their sum is 142. Can you find the numbers?
Answer by checkley77(12844)

x-y=38 or x=38+y
x+y=142
Now replace x with (38+y) in the second equation & solve for y.
(38+y)+y=142
38+2y=142
2y=142-38
2y=104
y=104/2
y=52 ans.
x-52=38
x=38+52
x=90 ans.
Proof:
52+90=142
142=142